3.53 The time from when a patient is discharged from North Shore Hospital to the time the discharged patient’s bed is ready to be assigned to a new patient is referred to as the bed assignment turnaround time. If the bed turnaround time is excessive, it can cause problems with patient flow and delay medical procedures throughout the hospital. This can cause long waiting times for physicians and patients, thus creating customer dissatisfaction. The admissions RN has assigned a patient care associate to measure the bed turnaround time for a randomly selected bed each morning, afternoon, and evening for 30 days. Following are the bed turnaround time sample observations:

DAY BED TURNAROUND TIMES (MIN)
  1 127 135 167
  2 140 155 122
  3 112 128 97
  4 223 135 154
  5 181 155 160
  6 103 158 145
  7 146 135 167
  8 104 122 115
  9 136 158 137
10 145 163 106
11 84 146 125
12 169 152 208
13 216 124 163
14 190 178 103
15 148 205 144
16 157 151 126
17 142 102 95
18 166 178 159
19 177 211 204
20 98 91 158
21 133 160 152
22 212 131 138
23 180 165 134
24 88 126 108
25 95 156 138
26 156 202 177
27 144 157 165
28 184 171 106
29 138 142 155
30 150 99 148

a.Develop an x-chart to be used in conjunction with an R-chart using 3σ limits to monitor the bed turnaround times and indicate if the process is in control using these charts.

b.Is the hospital capable of consistently achieving bed turnaround times of 120 minutes ±15 minutes without improving the process?
Answer #1

(a)

Day x1 x2 x3 X-bar Range
1 127 135 167 143.0 40
2 140 155 122 139.0 33
3 112 128 97 112.3 31
4 223 135 154 170.7 88
5 181 155 160 165.3 26
6 103 158 145 135.3 55
7 146 135 167 149.3 32
8 104 122 115 113.7 18
9 136 158 137 143.7 22
10 145 163 106 138.0 57
11 84 146 125 118.3 62
12 169 152 208 176.3 56
13 216 124 163 167.7 92
14 190 178 103 157.0 87
15 148 205 144 165.7 61
16 157 151 126 144.7 31
17 142 102 95 113.0 47
18 166 178 159 167.7 19
19 177 211 204 197.3 34
20 98 91 158 115.7 67
21 133 160 152 148.3 27
22 212 131 138 160.3 81
23 180 165 134 159.7 46
24 88 126 108 107.3 38
25 95 156 138 129.7 61
26 156 202 177 178.3 46
27 144 157 165 155.3 21
28 184 171 106 153.7 78
29 138 142 155 145.0 17
30 150 99 148 132.3 51
Avg. 146.79 47.47
X_double-bar R-bar

Note the following coefficients from a standard 3-sigma X-bar and R table for sample size n = 3.

A2 = 1.023
D3 = 0
D4 = 2.574
d2 = 1.693

UCLX = X_double-bar + A2 * R-bar = 146.79 + 1.023 * 47.47 = 195.35
LCLX = X_double-bar - A2 * R-bar = 146.79 - 1.023 * 47.47 = 98.23

UCLR = D4 * R-bar = 2.574 * 47.47 = 122.18
LCLR = D3 * R-bar = 0 * 47.47 = 0

The process has shown out-of-control behavior in Day-19 where the sample mean has crossed the UCLX.

(b)

We following the above process in re-constructing the control limits by removing the Day-19 data assuming that it is due to a special cause variation that has been resolved. The result is as follows:

In the modified chart, R-bar = 47.93 and X_double-bar = 145.05

We estimate the process mean (μ) and process stdev (σ) as follows:

μ = X_double-bar = 145.05
σ = R-bar / d2 = 47.93/1.693 = 28.31

USL = 120+15 = 135
LSL = 120 - 15 = 105

Cpu = (USL - μ) / 3.σ = (135 - 145.05) / (3*28.31) = -0.12
Cpl = (μ - LSL) / 3.σ = (145.05 - 105) / (3*28.31) = 0.47

Cpk = Min(Cpu, Cpl) = -0.12 which is less than 1.0. So, the process is incapable of producing withing the specificaion for a 3-sigma standard.